OpenStudy (anonymous):

Stuck on an algebra 2 question

OpenStudy (anonymous):

sqrt(x - 8) + 3 = 1

OpenStudy (anonymous):

Is the 3 under the square root?

OpenStudy (anonymous):

move the 3 to the other side of the equation, then square the problem

OpenStudy (anonymous):

I know that there is no solution, but I'm getting stuck here: sqrt(x - 8) + 3 = 1 sqrt(x - 8) = -2 I'm not sure if when I square both sides what the "-2" becomes. No, the 3 isn't.

OpenStudy (anonymous):

The -2 becomes a 4 when squared

OpenStudy (anonymous):

Is this right? sqrt(x - 8) + 3 = 1 sqrt(x - 8) = -2 x - 8 = 4 x = 4 + 8 x = 12

OpenStudy (anonymous):

Yes that is :)

OpenStudy (kc_kennylau):

no because you only take the POSITIVE value when evaluating square roots

OpenStudy (kc_kennylau):

therefore it has no solution

OpenStudy (anonymous):

Thank you, guys! Can you help me check it or whatever?

OpenStudy (anonymous):

Yeah I know there is no solution, now I have to check it to show that it's extraneous.

OpenStudy (kc_kennylau):

come on only me insist that there is no solution doesn't mean that i'm wrong

OpenStudy (anonymous):

It does have a solution. sqrt (12-8)+3=1 sqrt(4)+3=1 2+3=1 5=1 It's extraneous

OpenStudy (kc_kennylau):

when you do square root, you only take the POSITIVE value

OpenStudy (kc_kennylau):

as zeinas said sqrt(12-8)+3=sqrt4+3=2+3=5 not 1

OpenStudy (kc_kennylau):

If x^2=4 then x can be +2 or -2 but sqrt4 is just +2

OpenStudy (anonymous):

He's right, the x=12 is extraneous so it's not a solution.

OpenStudy (anonymous):

KC is right.

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Square_root "Although the principal square root of a positive number is only one of its two square roots, the designation "the square root" is often used to refer to the principal square root. For positive a, the principal square root can also be written in exponent notation, as \(a^{\frac12}\)."

OpenStudy (anonymous):

Kc is right: sqrt(x - 8) + 3 = 1 sqrt(x - 8) = -2 x - 8 = 4 x = 4 + 8 x = 12 Check sqrt(12 - 8) + 3 = 1 12 - 8 + 3 = -1 4 + 3 = 1 7 = 1 The solution is extraneous.

OpenStudy (anonymous):

Thanks for reminding me of the "rules". the square root of something can never be negative. Wolfram reinforces.

OpenStudy (anonymous):

P.S I THINK I did that right haha

OpenStudy (anonymous):

lol You should have done the 12-8 first, which is 4, then find the sqrt of 4 which is 2, then add that to three, but you were right about it being extraneous.

OpenStudy (anonymous):

Oh okay. Can you guys explain one more thing to me? Please :) When you're checking it, how does it go from = 1 = -1 = 1 = 1

OpenStudy (kc_kennylau):

Here would be my solution: Consider the equation \(\sqrt{x-8}+3=1\). \[\begin{array}{rcl}\sqrt{x-8}+3&=&1\\\sqrt{x-8}&=&1-3=-2\end{array}\]Since a square root of a number cannot be negative, it has no solution.

OpenStudy (anonymous):

@HHS6281019 It's not supposed to change to -1. It should be +1 all the way. It was probably a typo.

OpenStudy (anonymous):

Okay lol, that really confused me. Thanks again, guys! I'm not sure who to give a medal to?

OpenStudy (anonymous):

give it to me

OpenStudy (kc_kennylau):

Check for x=12: \[\begin{array}{}LHS&=\sqrt{12-8}+3\\&=\sqrt4+3\\&=2+3\\&=5\\\\RHS&=1\end{array}\] Therefore x=12 is not a solution. (P.S. I've already proved that it has no solution, so this step is unnecessary)

OpenStudy (anonymous):

id say KC

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

You don't have to give a medal. We all helped in a way

OpenStudy (kc_kennylau):

@zeinas. yep :D

OpenStudy (anonymous):

psh i derserve that medal

OpenStudy (kc_kennylau):

here's your medal :)

OpenStudy (anonymous):

lol kc u actually gave me one?

OpenStudy (anonymous):

there ya go

OpenStudy (kc_kennylau):

lolz

OpenStudy (anonymous):

i think there is a solution in complex numbers which is x =4i+8

OpenStudy (kc_kennylau):

remember \(\Large\sqrt i=\dfrac1{\sqrt2}(1+i)\)

OpenStudy (kc_kennylau):

Check: sqrt(4i+8-8)+3 =sqrt(4i)+3 =2sqrt(i)+3 =2/sqrt(2)+2i/sqrt(2)+3 =sqrt(2)+sqrt(2)i+3

OpenStudy (kc_kennylau):

@mark_boltzman therefore x=4i+8 is not a solution :)

OpenStudy (anonymous):

you r right i did a mistake :)

OpenStudy (kc_kennylau):

I also thought about complex number at the first place so it's fine :)

OpenStudy (anonymous):

Wow that's some complicated stuff haha

OpenStudy (kc_kennylau):

lolz :P